OpenStudy (anonymous):

You downloaded a video game to your computer. You have a 60-minute free trial of the game. It takes 5 minutes to set up the game, and 7 minutes to play each level. You are trying to find out how many levels, L, you can play. You can play whole levels or fractional parts of levels. The inequality representing this situation is: 5+7L≤60 What is the solution set for L, the number of levels you can play in the 60-minute time limit?

OpenStudy (ja1):

For this I suggest forgetting about the fact that you have an inequality and simply think of it as an equation: \[5 + 7l \le 60\]subtract 5 from both sides\[7l \le 55\]Can you go from there?

OpenStudy (anonymous):

7.85

OpenStudy (ja1):

Yupp! Perfect.

OpenStudy (ja1):

Let's plug it in to make sure it's correct. \[5 + 7(7.85) \le 60\]\[5 + 54.95 \le 60\]\[59.95 \le 60\]Is that correct?

OpenStudy (anonymous):

yes because it is just under 60

OpenStudy (ja1):

Oh wait, I thought @arrdasha was asking the question. Please refrain from giving out answers in the future arrdasha.

OpenStudy (ja1):

@sisi_swagger are you here?

OpenStudy (anonymous):

ya

OpenStudy (ja1):

Do you understand the problem?

OpenStudy (anonymous):

The answer is wrong

OpenStudy (anonymous):

The solution set of the inequality is L≤ 55 7

OpenStudy (anonymous):

The answer is L is greater than or equal to 55/7

OpenStudy (ja1):

Why did you switch the Less Than for a Greater Than symbol? You only do that if you divide by a negative number.

OpenStudy (anonymous):

idk the website that asked the question said that was the right answer

OpenStudy (ja1):

That's weird, perhaps I did something wrong?